# Quantum Monte Carlo and Stabilizer States

### v1

@@ III.3 Simulation results
We also performed an exact diagonalization calculation using the truncated partition function of eq. (50) and the full partition function ($L=\infty$). The results are shown in Fig. 2. The Monte Carlo results are in good agreement with the exact results, the relative error being smaller than $1\%$ , see inset. The author’s C++ implementation can be found in the GitHub repository Jeevanesan .
@@
IV Stabilizer updates for the transverse field Ising model and $\mathbb{Z}_{2}$ gauge theories
@@
The stabilizer scheme that we have described is not limited to the toy model above. Instead, a large number of spin-$1/2$ models can be reformulated in this way. As an example, the transverse field Ising model reads
$H$ $=$ $-\sum_{\langle ij\rangle}\frac{Z_{i}Z_{j}+\mathbf{1}}{2}-h\sum_{i}\frac{X_{i}+\mathbf{1}}{2}$ (51)
$=$ $-\sum_{\langle ij\rangle}\tilde{\Pi}_{ij}-h\sum_{i}\Pi_{i}$ (52)
@@
where we have added constants to the operators to turn them into projectors. The second term is the $\Pi_{i}$ operator that we have already encountered, while the first term is a new kind of projector that we denote by $\tilde{\Pi}_{ij}$. Above, we have given the update rules when $\Pi_{i}$ acts. When $\tilde{\Pi}_{ij}$ acts, the rules are nearly as simple. Again, we retain all the generators that commute with $Z_{i}Z_{j}$. We list all operators that anti-commute with $Z_{i}Z_{j}$, replace the first of these by $Z_{i}Z_{j}$ and make new generators by taking products as before. We also need to update the overall factor ${F}$ to account for potential changes of the norm of a state. Since $\Pi_{i}$ and $\tilde{\Pi}_{ij}$ have all non-negative entries in the $Z$-basis, the matrix elements $\mathcal{M}$ are all non-negative for $h>0$, thereby avoiding the sign problem.
@@
In $\mathbb{Z}_{2}$ lattice gauge theories it is usual to have products of multiple spin operators appearing in the Hamiltonian. An example is Kitaev’s toric-code Hamiltonian Kitaev (2003), which contains products of four spin-1/2 operators as star and plaquette terms. We note that $\Pi_{1}\equiv{(\mathbf{1}+X_{1}X_{2}X_{3}X_{4})}/2$ and $\Pi_{2}\equiv{(\mathbf{1}+Z_{1}Z_{2}Z_{3}Z_{4})}/2$ are projection operators that are non-negative in the $Z$ basis. Their update rules are similar to the ones for $\tilde{\Pi}_{ij}$. In this way, the partition functions of certain $\mathbb{Z}_{2}$ lattice gauge theories will be simulable without sign problems.
@@ V Conclusion
To summarize, we have demonstrated a scheme to remove the no-branching condition in the SSE quantum Monte-Carlo algorithm. This allowed us to straightforwardly deal with the CNOT toy model. We also indicated how the simulation would proceed in the case of the transverse field Ising model and certain $\mathbb{Z}_{2}$ gauge theories involving multi-spin operators.
There are several interesting directions to pursue. The simulation of Clifford circuits in this paper, although requiring only polynomial time and space complexity, was performed in the simplest possible way. Therefore, it seems likely that more sophisticated approaches, like the graph state simulation of Anders and Briegel (2006), will provide significant speedups.
In terms of applications, it is worth mentioning that lattice gauge theories, once the domain of high-energy physics Creutz (1983), have recently garnered attention from the cold atoms community as candidates to test quantum simulators Aidelsburger et al. (2022). Another natural context in which lattice gauge theories appear is the study of spin liquids Savary and Balents (2016). It is worth exploring whether the use of stabilizer states could unlock the SSE technique for these models, especially in cases where an efficient updating scheme is difficult to design or unavailable.

### v2

@@ III.3 Simulation results
We also performed an exact diagonalization calculation using the truncated partition function of eq. (51) and the full partition function ($L=\infty$), see App. A for details. The data is shown in Fig. 2. The Monte Carlo results are in good agreement with the exact results, the relative error being smaller than $1\%$ , see inset. The author’s C++ implementation of the Monte Carlo algorithm and the Python implementation of the exact diagonalization can be found in the GitHub repository Jeevanesan .
@@
IV Stabilizer updates for the transverse-field Ising model and $\mathbb{Z}_{2}$ gauge theories
@@
The stabilizer scheme that we have described is not limited to the toy model above. Instead, a large number of spin-$1/2$ models can be reformulated in this way. As another example, we apply the technique to the transverse-field Ising model defined by
$H$ $=$ $-J\sum_{\langle ij\rangle}\frac{Z_{i}Z_{j}+\mathbf{1}}{2}-h\sum_{i}\frac{X_{i}+\mathbf{1}}{2}$ (52)
$=$ $-J\sum_{\langle ij\rangle}\tilde{\Pi}_{ij}-h\sum_{i}\Pi_{i}.$ (53)
@@
We have added constants to the operators to turn them into projectors. The second term is the $\Pi_{i}$ operator that we have already encountered, while the first term is a new kind of projector that we denote by $\tilde{\Pi}_{ij}$. Above, we have given the update rules when $\Pi_{i}$ acts. When $\tilde{\Pi}_{ij}$ acts, the rules are nearly as simple. Firstly, retain all the generators that commute with $Z_{i}Z_{j}$. Then list all operators that anti-commute with $Z_{i}Z_{j}$, replace the first of these by $Z_{i}Z_{j}$ and make new generators by taking products as before. Since $\Pi_{i}$ and $\tilde{\Pi}_{ij}$ are projection operators, we need to keep track of the overall factor ${F}$ to account for potential changes of the norm of a state. The update rule is the same as before.
Since $\Pi_{i}$ and $\tilde{\Pi}_{ij}$ have all non-negative entries in the $Z$-basis, the matrix elements $\mathcal{M}$ are all non-negative for $h>0$, thereby avoiding the sign problem. The author’s stabilizer Monte Carlo code for the transverse-field Ising model is available in the GitHub repository Jeevanesan together with the corresponding exact diagonalization code that was used to compare the numerics. As an example, we show in Fig. 3 the average thermal energy of the system as a function of temperature. To compare with exact diagonalization, we have used $N=10$ Ising spins. The system is again thermalized for $0.5\times 10^{5}$ cycles and measurements are taken in the next $0.5\times 10^{5}$ cycles. The Monte Carlo data (dots) agree well with the exact diagonalization results, the relative error being less than $1\%$, as shown in the inset of Fig. 3.
Figure 3: Thermal average energy of the transverse-field Ising model, defined in eq. (53). We simulate the system for $N=10$ qubits and an external field of $h/J=3.0$ for various values of the expansion order $L$. The continuous lines are exact results obtained by calculating the mean energy with the truncated partition function in eq. (51) using exact diagonalization. The points are the results of the quantum Monte Carlo simulation. The relative error at each data point is always less than $1\%$, an example for $L=40$ is shown in the inset.
@@
Finally, we note that an extension of the presented approach to $\mathbb{Z}_{2}$ lattice gauge theories is straightforward. In such models it is usual to have products of multiple spin operators appearing in the Hamiltonian. An example is Kitaev’s toric-code Hamiltonian Kitaev (2003), which contains products of four spin-1/2 operators as star and plaquette terms. Such combinations of operators can be rewritten in terms of projection operators $\Pi_{1}\equiv{(\mathbf{1}+X_{1}X_{2}X_{3}X_{4})}/2$ and $\Pi_{2}\equiv{(\mathbf{1}+Z_{1}Z_{2}Z_{3}Z_{4})}/2$ that are non-negative in the $Z$ basis. Their update rules are similar to the ones for $\tilde{\Pi}_{ij}$. In this way, the thermodynamics of certain $\mathbb{Z}_{2}$ lattice gauge theories will be simulable without sign problems.
@@ V Conclusion
To summarize, we have demonstrated a scheme to remove the no-branching condition in the SSE quantum Monte Carlo algorithm by utilizing Clifford stabilizer states. In the language of word-line trajectories, our algorithm evaluates multiple branches of the wave function simultaneously, instead of averaging over trajectories one by one. This allowed us to straightforwardly deal with the CNOT toy model. Next, it was shown that the simulation of the transverse-field Ising model is similarly effortless, requiring only minor modifications of the technique. Finally, we indicated how the approach may be also leveraged in the simulation of $\mathbb{Z}_{2}$ gauge theories that usually involve terms of multi-spin operators. The present paper focused on thermal averages, but the SSE technique can also be adapted to compute ground state properties Sandvik (2010). The stabilizer approach illustrated here readily extends to these calculations.
There are several interesting directions to pursue. The simulation Jeevanesan of Clifford circuits in this paper, although requiring only polynomial time and space complexity, was performed in the simplest possible way. Therefore, it seems likely that more sophisticated approaches, like the tableau-based simulation of Aaronson and Gottesman (2004) or the graph-state simulation of Anders and Briegel (2006), will provide significant speedups.
In terms of applications, it is worth mentioning that lattice gauge theories, once the domain of high-energy physics Creutz (1983), have recently garnered increasing attention from the cold atoms community as candidates to investigate quantum simulators Wiese (2013); Bauer et al. (2023); Aidelsburger et al. (2022); Banuls et al. (2020). It is worth exploring whether stabilizer states could unlock the SSE technique for these models, especially in cases where an efficient updating scheme is difficult to design or otherwise unavailable. Although this approach does not mitigate the sign problem generally, the class of simulable models might nevertheless be expanded. In this way, the stabilizer-SSE algorithm may provide a generic tool to simulate models and make it unnecessary to tailor update algorithms specifically to fit a particular model.
Finally, we note that Hamiltonians of interacting qubits with controlled-X and controlled-Z operations, somewhat resembling eq. (10), have been explored in the literature as models for gapped phases of matter Chen et al. (2011); Stephen and Chen (2024). This is a further potential playground for the stabilizer-SSE approach.
VI Acknowledgement
It is a pleasure to thank Pok Man Tam and Pak Kau Lim for an interesting afternoon of discussion.
Appendix A Details on the Exact Diagonalization Calculation
In the main text, we gauged the quality of the Monte Carlo algorithm by comparing the average thermal energy to the exact diagonalization (ED) results. The ED implementation is simple and can be found in the author’s GitHub repository Jeevanesan . The code uses NumPy Harris et al. (2020) to construct the Hamiltonian matrix. For $N$ qubits, the latter has dimension $2^{N}\times 2^{N}$. The implementation starts by constructing the $\text{CX}_{i}$ and $\Pi_{i}$ operators for each site $i$ via Kronecker products of Pauli matrices and identity operators. This yields the Hamiltonian $H$ in matrix form and eigenvalues $E_{i}$ are straightforwardly computed. As explained in the main text, we construct the mean energy
$\langle H\rangle=-\partial_{\beta}\log Z(\beta)$ (54)
from the partition function $Z=-\text{Tr}\left[\exp({-\beta H}\right])$. We use the symbolic manipulation package SymPy Meurer et al. (2017) to compute the partition function
$Z(\beta)=\sum_{i=1}^{2^{N}}e^{-\beta E_{i}}$ (55)
using the known eigenvalues $E_{i}$. From this, we compute the average energy $\langle H\rangle=-\partial_{\beta}\log Z(\beta)$ by symbolical differentiation w.r.t. $\beta$ and thereby avoid ill-conditioned finite-difference calculations.
Similarly, we are also interested in the truncated partition function
We tabulate the values of $\text{Tr}\left[H^{n}\right]$ and use SymPy to construct the polynomials $Z_{L}(\beta)$ that we again differentiate, according to $\langle H\rangle_{L}=-\partial_{\beta}\log Z_{L}(\beta)$, to obtain the mean energy. Fig. 2 in the main text shows the resulting curves for $L=10,20,30,40$ and infinite order.
Our Hamiltonian matrix has dimension $D=2^{N}$. Since a general $D\times D$ matrix requires $O(D^{3})$ steps for the diagonalization and there are $2^{N}$ basis states, the computational complexity of the full diagonalization increases exponentially as $O(8^{N})$.